Question 6 1 pts A quality control lab has just purchased a new analytical instrument to replace the old one that has served for more than 10 years. Daily operation of the old instrument include running a sample of the same batch for 10 years as quality assurance measure. The old method yielded over all these years an average value of 10.5 ppm with a standard deviation of 0.2 ppm. The new instrument has been tested for a week and the first results are shown below 10.0 ppm 10.4 ppm 10.2 ppm 10.8 ppm 10.1 ppm 10.5 ppm 10.1 ppm Compare the old and new instrument, can you assume at 99% confidence level that the new instrument gives a different result than the old instrument? Perform a significance test and select the correct answer O texp 1.89 o 3.707 exp? tcnt, therefore reject null hypothesis. results can be considered different at 99% confidence interval. exp 1.89, tom 3.707 texpC tent, therefore retain null hypothesis. results cannot be considered different at 99% confidence interval O texp 2.04 to 3.707 expS tent, therefore retain null hypothesis: results cannot be considered different at 99% confidence interval texp 1.89; tet 3.499 texpC tent, therefore retain null hypothesis: results cannot be considered different at 99% confidence interval
Answer #1

The following data is given: S. No. (in Results ppm) 10.0 10.4 10.2 10.8 10.1 10.5 10.1 Determine the mean and standard devia

Determine the degree of freedom as shown below: Degree of freedom = 7-1 = 7-1 =6 This value is determined by examining the st